# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v2

@@ 1.3 Methodology: Homomorphic Cryptanalysis
The fundamental approach to achieving verifiability in these lightweight schemes is to introduce an additional secret value alongside the secret key of HE. Both vADDG and REP adopt an index set $S$ as a secret value, which determines the slot or ciphertext position. Similarly, in the case of PE, the element $\alpha$ from $\mathbb{Z}_{t}^{\times}$ serves as the secret value. If an adversary obtains knowledge of these secret values, i.e., $S$ and $\alpha$, a forgery attack becomes trivial by manipulating the secret encryption. Therefore, these schemes conceal the secret values within ciphertexts.
Our approach to attack the schemes is to obtain the secret values encrypted states and use them to generate a cheating circuit. In the case of vADDG and REP scheme, the core part of the attack is to check if two ciphertexts encrypt the same plaintext in encrypted states. It is used to generate an encryption of 1 in the position of ciphertexts with identical values through homomorphic operations. Similarly, in the case of PE, knowing $\alpha$ can break the scheme since verification is done by evaluating $Y$ at $\alpha$. In the current attack, however, some re-randomized procedure is adopted to prevent the computation of $\alpha$ and our attacks detour this obstacle by considering a special circuit from interpolation at $Y=0$ and $Y=\alpha$.
In any case, even without knowing the secret value itself, the adversary can homomorphically implement a cheating circuit by leveraging the functionalities of homomorphic encryption. If the cheating circuit cannot be evaluated within this homomorphic capacity, the attack fails. The vADDG scheme relies on this property; however, our attack circumvents this limitation by constructing a depth-1 cheating circuit.
@@ 1.4 Related Works
There are two major security concerns in information security: integrity and confidentiality. As data utilization expands beyond mere storage to computational applications, these concerns have evolved into Verifiable Computation (VC) and Homomorphic Encryption (HE), respectively. To achieve VC, various cryptographic primitives have been proposed, including SNARKs [LHW+25], Homomorphic MACs [CF13, GW13], and Homomorphic Signatures [CCRS24]. Meanwhile, since Gentry’s groundbreaking discovery [Gen09], HE has seen continuous advancements in both functionality and efficiency. Notable schemes such as BGV, BFV, TFHE, and CKKS [BGV12, FV12, CGGI20, CKKS17] have significantly improved HE’s practical applicability.
A natural follow-up question is whether VC and HE can be integrated, allowing both concerns to be addressed simultaneously. The client would be able to ensure not only the privacy of their data but also the integrity of the computation results. However, integrating verifiability into HE is not merely about verifying the correctness of homomorphic computations. Due to the inherent malleability of FHE ciphertexts, an active adversary—if integrity measures are absent—can manipulate ciphertexts to intentionally trigger incorrect decryption, leading to the IND-CPAD attack and undermining confidentiality [LM21, CCP+24]. Therefore, ensuring the integrity of homomorphic computations is not just an extension of verifiability but a critical requirement for achieving malicious security.
However, adding verifiability to HE is a non-trivial task. It is not even clear what the precise definition of VFHE should be. Although there exists research on maliciously secure FHE that proposes new security notions for ensuring integrity [MN24, BSW11], a well-established and standardized definition has yet to be clearly formulated. In practice, there are two major ways to construct Verifiable HE. One way is to apply integrity mechanisms to the ciphertext operations of HE. However, conventional integrity mechanisms such as SNARKs conflict with HE’s non-arithmetic ciphertext operations. [ABPS24, BCFK21, FGP14, FNP20, GNSV23]. Another research direction attempts to achieve verifiability by generating SNARK-based proof in encrypted computation [ACGaSV24, GGW24, GBK+24]. The state-of-the-art study in this field [GBK+24] claims that its approach is four orders of magnitude faster than methods that generate proof for heavy HE operations. However, these approaches fail to prevent IND-CPAD attacks since integrity can only be verified after decryption.
@@
The verifiable homomorphic encryption scheme targeted in this paper follows a different approach from the two major existing methods. These prior approaches aim to ensure integrity by leveraging the confidentiality of HE and by modifying the message encoding structure of existing HE schemes [CKP+24, ADDG24, SXLS24]. However, there is no well-established hardness assumption or security proof that justifies a reduction from the confidentiality of HE to its integrity, making the connection between the two notions unclear. Thus, these approaches are constrained by weaker adversary models: [CKP+24] assumes a covert adversary model, while [ADDG24] assumes an adversary capable of performing bootstrapping only once.
@@
Our attack leverages the homomorphic properties of HE to forge the underlying encoding structure. This idea was implicitly suggested in [CHLR18], and later formalized in the construction of vADDG [ADDG24], where the authors proposed a naive hardness assumption: that constructing such a cheating circuit within shallow depth is infeasible. This suggests the need for a deeper exploration of what kinds of computations are possible in various homomorphic encryption settings.
@@ A Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
  • $\textsf{ct}{}_{y}\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$ where $f\in\mathcal{F}.$
  • $y\leftarrow\textsf{Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y}).$
@@
Let $\Pi$ be a Homomorphic Encryption scheme. $\Pi$ is $\delta$-Correct if the evaluation on the ciphertext decrypts to the correct result.
$\Pr\left[\textsf{Dec}_{\sf sk}(\textsf{ct}{}_{y})=f(x):\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F})\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{\textsf{pk}{}}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta.$
@@
TFHE [CGGI20] is a FHE scheme that supports gate-level bootstrapping, where the (v)ADDG scheme is instantiated with. Its plaintext space is $\mathbb{Z}_{Q}$ and the message space $\mathbb{Z}_{P}$ is encoded into $\mathbb{Z}_{Q}$ for $P<Q$. Thereafter a plaintext is encrypted into the ciphertext space $\mathcal{C}=\mathbb{Z}_{Q}^{n+1}$ where $n$ is determined by the security parameter.
@@
BFV is an RLWE-based HE scheme that supports the computation of vectors of integers [Bra12, FV12], and it is the scheme in which Veritas is instantiated. Its message space is $\mathbb{Z}_{t}^{N}$, where $t=p^{r}$ is a power of a prime $p$. The ciphertext space is $\mathcal{R}^{2}_{q}=\mathbb{Z}_{q}[X]/(\Phi_{M}(X))$ for $q>t$, where $\Phi_{M}(X)\in\mathbb{Z}[X]$ is the $M$-th cyclotomic polynomial with degree $N=\varphi(M)$. Its multiplication is the Hadamard multiplication. One of the notable features of this scheme is that it supports rotation, $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which is the shifting action on $\mathbb{Z}_{t}^{N}$ with the aid of the rotation key $\textsf{rtk}{}_{i}$. Also, the multiplication between two BFV ciphertexts relies on the relinearization key rlk. As far as known, BFV bootstrapping requires the evaluation keys $\textsf{rtk}{}_{1}$ and rlk.
The BFV scheme is a RLWE based homomorphic encryption scheme that enables computations over vectors of integers. It serves as the underlying encryption scheme for Veritas. The message space of BFV is $\mathbb{Z}_{t}^{N}$, where $t=p^{r}$ denotes a prime power. Its ciphertext space is defined as $\mathcal{R}^{2}_{q}=\mathbb{Z}_{q}[X]/(\Phi_{M}(X))$, where $\Phi_{M}(X)\in\mathbb{Z}[X]$ is the $M$-th cyclotomic polynomial of degree $N=\varphi(M)$, and $q>t$ is a ciphertext modulus.
Multiplication in BFV is performed component-wise using Hadamard multiplication. A notable feature of BFV is its support for rotation operations $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which cyclically shift the message vector and require a corresponding rotation key $\textsf{rtk}{}_{i}$. Furthermore, ciphertext-ciphertext multiplications are enabled through the use of a relinearization key rlk. As of current knowledge, bootstrapping in BFV necessitates both the rotation keys $\textsf{rtk}{}_{1}$ and the relinearization key rlk. For further technical details, refer to [GV23].
@@ A Verifiable Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
  • $\textsf{ct}{}_{y}\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$ where $f\in\mathcal{F}$.
  • $b\leftarrow{\sf Verify_{\textsf{sk}}{}}(\textsf{ct}{}_{y})$, the client accepts if $b=0$ and rejects if $b=1$.
  • $y\leftarrow\textsf{Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y}).$
@@
A Verifiable Homomorphic Encryption scheme $\Pi$ is $\delta$-complete if the client always accepts a correct output.
$\Pr\left[0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y}):\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta$
@@
We now formally define the security of VHE. We assume that the underlying scheme $\Pi$ satisfies both $\delta$-correctness and $\delta$-completeness for some negligible $\delta$. Throughout this work, we consider two types of adversaries: malicious and covert.
@@
A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of malicious adversaries if the client rejects an incorrect output with overwhelming probability in $\lambda$ for any malicious PPT adversary $\mathcal{A}_{\sf mal}$.
$\Pr\left[\begin{aligned} 0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y})\\ \textsf{Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y})\neq f(x)\end{aligned}:\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\mathcal{A}^{\textsf{evk}}_{\sf mal}(\textsf{ct}{}_{x})\\ \end{aligned}\right]\text{ is negligible.}$.
@@
Originally introduced in the context of secure multiparty computation in [AL07], the covert adversary model can be extended to the VHE setting, given that HE can be regarded as a special case of two-party computation, and considered in [CKP+24]. The $\varepsilon$-covert adversary model, or simply the covert adversary model, assumes that any adversarial deviation is detected with probability at least $\varepsilon$. A customary choice of the deterrence factor $\varepsilon$ is a noticeable probability ($1/{\sf poly}(\lambda)$).Unless otherwise stated, we use the term covert adversary to refer to a $1/{\sf poly}(\lambda)$-covert adversary. In this paper, we provide an informal definition of the covert adversary model. For a formal treatment, we refer the reader to the Definition 3.4 from [AL07].
@@
Let $\varepsilon\in[0,1]$. A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of $\varepsilon$-covert adversaries if for any PPT adversary $\mathcal{A}$, any deviation from the prescribed protocol is detected with probability at least $\varepsilon$.
@@ 3 Cryptanalysis on vADDG
In [ADDG24], Albrecht et al. presented an Oblivious PRF candidate and its extension to a verifiable OPRF, known as the ADDG and vADDG schemes, respectively. The ADDG scheme utilizes TFHE to construct an OPRF candidate: the client outsources homomorphic encryption of a bit string $\mathbf{x}$, and the server homomorphically computes $\bf z=F(\mathbf{k},\mathbf{x})$ with its PRF key $\bf k$ and returns it to the client. In this process, the OPRF protocol requires that the server does not learn either $\mathbf{x}$ or $\mathbf{z}$.
We briefly explain how the PRF $F(\mathbf{k},\mathbf{x})$ is homomorphically calculated in ADDG. For the TFHE ciphertext space $\mathcal{C}$, we denote by $[x]_{P}\in\mathcal{C}$ a TFHE encryption of $x\in\mathbb{Z}_{P}$. For $\mathbf{x}=(x_{1},\dots,x_{k})\in\mathbb{Z}_{P}^{k}$, we denote
$[\mathbf{x}]_{P}=([x_{1}]_{P},[x_{2}]_{P},\dots,[x_{k}]_{P})\in\mathcal{C}^{k}.$
@@
to obtain an encryption of $\mathbf{y}$ over $\mathbb{Z}_{3}^{256}$. Finally, again using only homomorphic additions and constant multiplications, the server computes an encryption of an 82-trit string $\mathbf{z}\in\mathbb{Z}_{3}^{82}$, which is the output of the homomorphic evaluation of the PRF. A notable feature of the (v)ADDG scheme is that it consumes only a bootstrapping depth of one: By the removal of the key-switching key, this scheme prevents further bootstrapping executions after one bootstrapping.
@@
Figure 1: Workflow of the Homomorphic PRF suggested in [ADDG24]. The superscripted $\mathsf{sk}$ denotes that the secret key is $\mathsf{sk}$.
@@
    =$\mathfrak{R}=\{(\mathbf{x}^{\star}_{k},\mathbf{z}^{\star}_{k})\}_{1\leq k\leq\kappa}\quad(\mathbf{z}^{\star}_{k}=F(\mathbf{k},\mathbf{x}^{\star}_{k}))$
@@
    where $(\mathbf{x}^{\star}_{k_{j}},\mathbf{z}^{\star}_{k_{j}})\overset{\$}{\leftarrow}\mathfrak{R}$ and $\gamma=\alpha\nu+\beta$.

### v3

@@ 1.3 Methodology: Homomorphic Cryptanalysis
The fundamental approach to achieving verifiability in these lightweight schemes is to hide precomputed values via additional secrets alongside the secret key of HE. Both vADDG and REP adopt an index set $S$ as a secret value, which determines the slot or ciphertext position. Similarly, in the case of PE, the element $\alpha$ from $\mathbb{Z}_{t}^{\times}$ serves as the secret value. If an adversary obtains knowledge of these secret values, i.e., $S$ and $\alpha$, a forgery attack becomes trivial by manipulating the secret encryption. Therefore, these schemes conceal the secret values within ciphertexts.
Our approach to attack the schemes is to obtain the secret values encrypted states and use them to generate a cheating circuit. For vADDG and REP, the central idea of the attack is to homomorphically test whether two slots encrypt the same message. This enables the adversary to produce an encryption of 1 at positions corresponding to equal plaintexts via homomorphic operations. In contrast, for PE, the security hinges on the secrecy of $\alpha$, as verification is carried out by evaluating $Y$ at $\alpha$. While the ReQuadratization employs a re-randomization procedure to obscure $\alpha$ and prevent its computation, our attacks bypass this defense by analyzing the encoding structure of PE and devising a tailored circuit based on interpolation at $Y=0$ and $Y=\alpha$.
In any case, even without knowing the secret value itself, the adversary can homomorphically implement a cheating circuit by leveraging the functionalities of homomorphic encryption.
@@ 1.4 Related Works
There are two major security concerns in information security: integrity and confidentiality. As data utilization expands beyond mere storage to computational applications, these concerns have evolved into the notions of Verifiable Computation and Homomorphic Encryption, respectively. To achieve VC, various cryptographic primitives have been proposed, including SNARKs [LHW+25], Homomorphic MACs [CF13, GW13], and Homomorphic Signatures [CCRS24]. Meanwhile, since Gentry’s groundbreaking discovery [Gen09], HE has seen continuous advancements in both functionality and efficiency. Notable schemes such as BGV, BFV, TFHE, and CKKS [BGV12, FV12, CGGI20, CKKS17] have significantly improved HE’s practical applicability.
A natural follow-up question is whether VC and HE can be integrated, allowing both concerns to be addressed simultaneously. There are two major approaches to constructing Verifiable HE [ZWL+25]. One approach is to apply integrity mechanisms to the ciphertext operations of HE [ABPS24, BCFK21, FGP14, FNP20, GNSV23]. Another approach is to achieve verifiability by generating SNARK-based proofs in the encrypted domain [ACGaSV24, GGW24, GBK+24]. However, adding verifiability to HE is a non-trivial task. Both approaches suffer from significant efficiency limitations. Compared to native HE computations, this overhead is larger by four to five orders of magnitude.
For the former, conventional integrity mechanisms such as SNARKs are incompatible with the non-arithmetic ciphertext operations of HE. As a result, these operations must be emulated by arithmetic operations, which incurs significant overhead when proving them. For the latter approach, the plaintext space must be suitable for generating proofs, which leads to inefficiency in parameter selection, and the process of homomorphically evaluating proofs introduces substantial overhead.
@@
The verifiable homomorphic encryption scheme targeted in this paper follows a lightweight approach different from the two major existing methods. The idea is to achieve verifiability by hiding verification secrets under the semantic security of HE and by altering the message encoding of the underlying scheme accordingly [CKP+24, ADDG24, SXLS24]. However, there is no well-established hardness assumption or security proof that justifies a reduction from the confidentiality of HE to its integrity, making the connection between the two notions unclear.
@@
Our attack leverages the homomorphic properties of HE to forge the underlying encoding structure. This idea was implicitly suggested in [CHLR18], and later formalized in the construction of vADDG [ADDG24], where the authors proposed a naive hardness assumption that constructing such a cheating circuit within shallow depth is infeasible. This suggests the need for a deeper exploration of what kinds of computations are possible in various homomorphic encryption settings.
@@ A Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
[…]
@@
Let $\Pi$ be a Homomorphic Encryption scheme. $\Pi$ is $\delta$-Correct if the evaluation on the ciphertext decrypts to the correct result. For all $f$ and $x$, it should be satisfied that
$\Pr\left[{\sf Dec}_{\sf sk}(\textsf{ct}{}_{y})=f(x):\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F})\\ \textsf{ct}{}_{x}&\leftarrow{\sf Enc}_{\mathsf{\textsf{pk}{}}}(x)\\ \textsf{ct}{}_{y}&\leftarrow{\sf Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta.$
@@
TFHE [CGGI20] is a FHE scheme that supports gate-level bootstrapping, that the (v)ADDG scheme is instantiated with. Its plaintext space is $\mathbb{Z}_{Q}$ and the message space $\mathbb{Z}_{P}$ is encoded into $\mathbb{Z}_{Q}$ for $P<Q$. Thereafter a plaintext is encrypted into the ciphertext space $\mathcal{C}=\mathbb{Z}_{Q}^{n+1}$ where $n$ is determined by the security parameter.
@@
BFV is an RLWE-based HE scheme that supports the computation of vectors of integers [Bra12, FV12], and it is the scheme in which Veritas is instantiated. Its message space is $\mathbb{Z}_{t}^{N}$, where $N$ is a power of $2$ and $t=p^{r}$ is a power of a prime $p$. For $q>t$, the ciphertext space is $\mathcal{R}^{2}_{q}=\bigl(\mathbb{Z}[X]/(q,X^{M}+1)\bigr)^{2}$ where $M$ a power of $2$ and $N=M/d$ where $d$ is a multiplicative order of $p\in\mathbb{Z}_{M}^{\times}$. Multiplication is given by the Hadamard product, i.e., element-wise multiplication.
A notable feature of BFV is that it supports rotations $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which shift vectors in $\mathbb{Z}_{t}^{N}$ with the aid of the rotation key $\textsf{rtk}{}_{i}$. In addition, ciphertext–ciphertext multiplication requires the relinearization key rlk. As far as known, BFV bootstrapping requires the evaluation keys $\textsf{rtk}{}_{1}$ and rlk. For further technical details, refer to Appendix 0.A.
@@ A Verifiable Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
[…]
@@
A Verifiable Homomorphic Encryption scheme $\Pi$ is $\delta$-complete if the client always accepts a correct output. For all $f$ and $x$, it should be satisfied that
$\Pr\left[0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y}):\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ \textsf{ct}{}_{x}&\leftarrow{\sf Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow{\sf Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta$
@@
We now formally define the security of VHE. We assume that the underlying scheme $\Pi$ satisfies both $\delta$-correctness and $\delta$-completeness for some negligible $\delta$. Throughout this work, we consider two types of adversaries: malicious and covert. The following definition for malicious adversaries is a modification of the corresponding definition in [VKH23]. Specifically, the modification is the removal of the tag, as noted above.
@@
A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of malicious adversaries if the client rejects an incorrect output with overwhelming probability in $\lambda$ for any malicious PPT adversary $\mathcal{A}_{\sf mal}$. For all $f$, it should be satisfied that
$\Pr\left[\begin{aligned} 0\leftarrow{\sf Verify}_{\sf sk}(\textsf{ct}{}_{y})\\ {\sf Dec}_{\textsf{sk}{}}(\textsf{ct}{}_{y})\neq f(x)\end{aligned}:\begin{aligned} \textsf{pk}{},\textsf{sk}{}&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F}))\\ x&\leftarrow\mathcal{A}^{\textsf{evk}}_{\sf mal}\\ \textsf{ct}{}_{x}&\leftarrow{\sf Enc}_{\mathsf{pk}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\mathcal{A}^{\textsf{evk}}_{\sf mal}(\textsf{ct}{}_{x})\\ \end{aligned}\right]\text{ is negligible.}$
@@
Originally introduced in the context of secure multiparty computation in [AL07], the covert adversary model can be extended to the VHE setting, given that HE can be regarded as a special case of two-party computation, and was adopted as the adversary model in [CKP+24]. The $\varepsilon$-covert adversary model, assumes that any adversarial deviation is detected with probability at least $\varepsilon$. A customary choice of the deterrence factor $\varepsilon$ is a noticeable probability, e.g., $1/{\sf poly}(\lambda)$. Unless otherwise stated, we use the term covert adversary to refer to a $1/{\sf poly}(\lambda)$-covert adversary. In this paper, we provide an informal definition of the covert adversary model. For a formal treatment, we refer the reader to Definition 3.4 from [AL07].
@@
Let $\varepsilon(\lambda)\in[0,1]$. A Verifiable Homomorphic Encryption scheme $\Pi$ is sound in the presence of $\varepsilon$-covert adversaries if for any PPT adversary $\mathcal{A}$, any deviation from the prescribed protocol is detected with probability at least $\varepsilon$.
@@ 3 Cryptanalysis of vADDG
In [ADDG24], Albrecht et al. presented an Oblivious PRF candidate and its extension to a verifiable OPRF, denoted as the ADDG and vADDG schemes, respectively. The ADDG scheme utilizes TFHE to construct an OPRF candidate: the client outsources homomorphic encryption of a bit string $\mathbf{x}$, and the server homomorphically computes $\mathbf{z}=F(\mathbf{k},\mathbf{x})$ with its PRF key $\bf k$ and returns it to the client.
We briefly explain how the PRF $F(\mathbf{k},\mathbf{x})$ is homomorphically calculated in ADDG. For the TFHE ciphertext space $\mathcal{C}$, we denote by $x]_{P}\in\mathcal{C}$ a TFHE encryption of $x\in\mathbb{Z}_{P}$. For $\mathbf{x}=(x_{1},\dots,x_{k})\in\mathbb{Z}_{P}^{k}$, we denote
$\mathbf{x}]_{P}=([x_{1}]_{P},[x_{2}]_{P},\dots,[x_{k}]_{P})\in\mathcal{C}^{k}.$
@@
to obtain an encryption of $\mathbf{y}$ over $\mathbb{Z}_{3}^{256}$. Finally, again using only homomorphic additions and constant multiplications, the server computes an encryption of an 82-trit string $\mathbf{z}\in\mathbb{Z}_{3}^{82}$, which is the output of the homomorphic evaluation of the PRF. A notable feature of the (v)ADDG scheme is that it consumes only a bootstrapping depth of one: By the removal of the key-switching key, this scheme prevents further bootstrapping executions after the $\sf CPPBS$.
@@
Figure 1: Workflow of the Homomorphic PRF suggested in [ADDG24]. The superscripted $\mathsf{sk}$ (resp. $\mathsf{sk^{\prime}}$) denotes that the secret key is $\mathsf{sk}$ (resp. $\mathsf{sk^{\prime}}$). $\bf k$ is the PRF key.
@@
    $\mathfrak{R}=\{(\mathbf{x}^{\star}_{k},\mathbf{z}^{\star}_{k})\}_{1\leq k\leq\kappa}\quad(\mathbf{z}^{\star}_{k}=F(\mathbf{k},\mathbf{x}^{\star}_{k}))$
@@
    where $(\mathbf{x}^{\star}_{k_{j}},\mathbf{z}^{\star}_{k_{j}})\overset{\mathdollar}{\leftarrow}\mathfrak{R}$ and $\gamma=\alpha\nu+\beta$.